Ghana’s inflation is forecasted to be 10.2% over the coming year whilst that of South

Africa is forecasted to be 6.5%. The current exchange rate between Ghana Cedi and the

South African Rand is 0.32 ZAR per 1 GHS.

a) How should we quote the exchange rate between Ghana Cedi and the South

African Rand (ZAR) in a year’s time to avoid arbitrage?

b) A Ghanaian company is importing goods worth ZAR 20m in a year’s time, how

much GHS will the company require to import the goods?

c) If the actual rate at the end of the year is 0.35 ZAR per 1GHS, what is the

absolute forecast error for the forecast in (a)?


1
Expert's answer
2021-08-03T12:37:10-0400

a)

forward rate = spot rate"\times" [(1+inflation in ghana) "\div" (1+inflation in South africa)]

"= 0.32\times[\frac{(1+10.2\%)} { (1+6.5\%)} ]"

= 0.33 ZAR per 1 GHS(rounded to two decimals)

b)

"GHS\space require =\frac{ 20 \space million }{ 0.33}"

= 60.61million GHS (rounded to two decimals)

c)

when actual rate = 0.35

"GHS \space required = \frac{20\space million }{ 0.35}=

= 57.14\space million"

absolute fore cast error (in percentage) = (actual value - expected value) "\div" actual value

"= \frac{(57.14 - 60.61) }{ 57.14}"

= 6.06%(rounded to two decimals)